# MeLA: A Metacognitive LLM-Driven Architecture for Automatic Heuristic Design

### v2

@@ Introduction
To break this impasse, the field of Automatic Heuristic Design (AHD) emerged, aiming to automate the creation of heuristics. While foundational AHD paradigms like Hyper-Heuristics (HHs) (Pillay and Qu 2018) and Genetic Programming (GP) (Burke et al. 2007) were a conceptual leap forward, their creative potential is fundamentally capped. By design, they are tethered to human guidance, operating within search spaces defined by expert-supplied components or primitives (Mei et al. 2022). This inherent limitation prevents them from achieving true design autonomy, leaving the grand challenge of fully automated discovery unsolved.
@@
1:Input: Problem $P$, Generation Limit $L$
@@
6:   $f(H),E\leftarrow\text{Evaluate}(H)$
7:   if $E\neq\emptyset$ then
8:     $H,Th\leftarrow\text{CorrectErrors}(H,E,Th)$
9:     $f(H)\leftarrow\text{Evaluate}(H)$
10:   end if
11:   $Meta\leftarrow\text{AnalyzePerformance}(Th,f(H),$
12:                  $E,Best(H))$
13:   $H,Th\leftarrow\text{GenerateNext}(Meta)$
14:  end for
15:  return $Best(H)$
16:end procedure
@@
The optimization loop begins with an initial problem Analyze, which informs the InitializeHeuristics step, where the first population of heuristics $H$ and their corresponding thought processes $Th$ are generated. Subsequently, the core iterative cycle commences. First, all heuristics in the current population are subjected to Evaluation, which yields their fitness scores $f(H)$ and records any execution errors $E$. If $E$ are detected, the CorrectErrors procedure is invoked to repair the faulty heuristics and update the population $H$, thought processes $Th$ and fitness $f(H)$. Following this, the AnalyzePerformance stage initiates a metacognitive review $Meta$. It assesses the recorded thought processes $Th$, the fitness landscape of the population $f(H)$, the errors $E$ and the characteristics of the best-performing heuristic $Best(H)$ to generate a new set of refined prompts. These prompts are then used by the GenerateNext function to produce the next generation of heuristics $H$ and thoughts $Th$. This cycle of evaluation, correction, analysis, and generation repeats, driving the continuous refinement of the underlying problem-solving strategies.
@@
  • Generation Prompts: These prompts create the heuristics. The Initial Generation prompt uses the problem analysis to produce a diverse population of $N$ heuristics. Subsequently, the Metacognitive Generation prompt uses insights from the metacognition step to evolve this population, generating $N$ refined heuristics for the next iteration.
@@
Have PA 221.50221.50221.50 70.00% 4434.144434.144434.14 100.00%
No PA 221.80221.80221.80 56.67% 4447.324447.324447.32 93.33%
Initial 219.61219.61219.61 86.67% 5335.475335.475335.47 95.00%
Meta-1 207.16207.16207.16 90.00% 4311.724311.724311.72 90.00%
Meta-2 207.20207.20207.20 90.00% 3642.883642.883642.88 90.00%
Meta-3 3530.783530.783530.78 100%
@@ A.5 Problems
$\bullet$ WSN: Wireless Sensor Networks deployment problem involving 200 fixed-position Sensing Nodes (SNs) and 50 Convergence Nodes (CNs) whose positions and transmission powers need to be optimized. The system aims to achieve three primary objectives: ensuring all SNs are connected to at least one CN, maintaining connectivity among all CNs, and minimizing total power consumption while meeting technical requirements. The network operates with CNs having a 20-unit communication range and a capacity constraint of serving no more than 15 SNs each, using a path loss model that includes a 55 dB base loss, 2.5 path loss exponent, additional quadrant-based loss factors ($\beta_{x}$, $\beta_{y}$), and a minimum required signal strength of -85 dBm. The optimization employs a population-based metaheuristic approach with 50 search agents running for 100 iterations, where each solution represents CN configurations through three parameters (x position, y position, and power level) bounded within [0,50] for coordinates and [0,30] dBm for power. The fitness function combines multiple components including coverage penalties (10$\times$ per uncovered SN), connectivity penalties (1000 for disconnected CN networks), power uniformity penalties (100$\times$ for excessive standard deviation beyond 1 dB), and the true fitness metric of total power consumption converted from dBm to milliwatts. A feasible solution must satisfy all constraints by covering all 200 SNs, maintaining full CN connectivity, keeping power uniformity within limits, and achieving a fitness score below the 1000 penalty threshold, with the framework providing convergence tracking and visualization capabilities to monitor optimization progress throughout this complex multi-constraint problem. For details of the parameters and constraints, please refer to Table A4 and Table A6.

### v3

@@ Introduction
To break this impasse, the field of Automatic Heuristic Design (AHD) emerged, aiming to automate the creation of heuristics. While foundational AHD paradigms like Hyper-Heuristics ($H$s) (Pillay and Qu 2018) and Genetic Programming (GP) (Burke et al. 2007) were a conceptual leap forward, their creative potential is fundamentally capped. By design, they are tethered to human guidance, operating within search spaces defined by expert-supplied components or primitives (Mei et al. 2022). This inherent limitation prevents them from achieving true design autonomy, leaving the grand challenge of fully automated discovery unsolved.
@@
1:Input: Problem $P$, Number of Initializations NN, Generation Limit $L$
@@
6:  end for
7:  for $j\leftarrow N+1$ to $L$ do
8:   $f(H),E\leftarrow\text{Evaluate}(H)$
9:   if $E\neq\emptyset$ then
10:     $H,Th\leftarrow\text{CorrectErrors}(H,E,Th)$
11:     $f(H)\leftarrow\text{Evaluate}(H)$
12:   end if
13:   $Meta\leftarrow\text{AnalyzePerformance}(Th,f(H),$
14:                  $E,Best(H))$
15:   $H,Th\leftarrow\text{GenerateNext}(Meta)$
16:  end for
17:  return $Best(H)$
18:end procedure
@@
The optimization loop begins with an initial problem Analyze, which informs the InitializeHeuristics step, where the first population of heuristics $H$ and their corresponding thought processes ThTh are generated. Subsequently, the core iterative cycle commences. First, all heuristics in the current population are subjected to Evaluation, which yields their fitness scores $f(H)$ and records any execution errors $E$. If $E$ are detected, the CorrectErrors procedure is invoked to repair the faulty heuristics and update the population $H$, thought processes ThTh and fitness $f(H)$. Following this, the AnalyzePerformance stage initiates a metacognitive review MetaMeta. It assesses the recorded thought processes ThTh, the fitness landscape of the population $f(H)$, the errors $E$ and the characteristics of the best-performing heuristic $Best(H)$ to generate a new set of refined prompts. These prompts are then used by the GenerateNext function to produce the next generation of heuristics $H$ and thoughts ThTh. This cycle of evaluation, correction, analysis, and generation repeats, driving the continuous refinement of the underlying problem-solving strategies.
@@
  • Generation Prompts: These prompts create the heuristics. The Initial Generation prompt uses the problem analysis to produce a diverse population of NN heuristics. Subsequently, the Metacognitive Generation prompt uses insights from the metacognition step to evolve this population, generating NN refined heuristics for the next iteration.
@@
Have PA $221.50$ 70.00% $4434.14$ 100.00%
No PA $221.80$ 56.67% $4447.32$ 93.33%
Initial $219.61$ 86.67% $5335.47$ 95.00%
Meta-1 $207.16$ 90.00% $4311.72$ 90.00%
Meta-2 $207.20$ 90.00% $3642.88$ 90.00%
Meta-3 $3530.78$ 100%
@@ A.5 Problems
[…]